# Cream: Consistency Regularized Self-Rewarding Language Models

### v2

@@ Abstract
Recent self-rewarding large language models (LLM) have successfully applied LLM-as-a-Judge to iteratively improve the alignment performance without the need of human annotations for preference data. These methods commonly utilize the same LLM to act as both the policy model (which generates responses) and the reward model (which scores and ranks those responses). The ranked responses are then used as preference pairs to train the LLM via direct alignment technologies (e.g. DPO). However, it is noteworthy that throughout this process, there is no guarantee of accuracy in the rewarding and ranking, which is critical for ensuring accurate rewards and high-quality preference data. Empirical results from relatively small LLMs (e.g., 7B parameters) also indicate that improvements from self-rewarding may diminish after several iterations in certain situations, which we hypothesize is due to accumulated bias in the reward system. This bias can lead to unreliable preference data for training the LLM. To address this issue, we first formulate and analyze the generalized iterative preference fine-tuning framework for self-rewarding language model. We then introduce the regularization to this generalized framework to mitigate the overconfident preference labeling in the self-rewarding process. Based on this theoretical insight, we propose a Consistency Regularized sElf-rewarding lAnguage Model (CREAM) that leverages the rewarding consistency across different iterations to regularize the self-rewarding training, helping the model to learn from more reliable preference data. With this explicit regularization, our empirical results demonstrate the superiority of CREAM in improving both reward consistency and alignment performance. The code is publicly available at https://github.com/Raibows/CREAM.
@@ 1 Introduction
Large language models (LLMs) have demonstrated impressive capabilities across a range of tasks, including natural language understanding and generation (Radford et al., 2019). However, LLMs also face alignment challenges, such as generating hallucinations and harmful outputs (Ji et al., 2023). To address these issues, a series of research efforts have explored preference learning methods like Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017) and direct alignment techniques such as Direct Preference Optimization (DPO) (Rafailov et al., 2023) to better align LLMs with human values and preferences. These alignment methods often rely heavily on large amounts of preference pairs, which are essential for both RLHF and direct alignment training. However, collecting human-annotated preference pairs is time-consuming and labor-intensive, posing significant challenges to the scalability and efficiency of these alignment approaches.
Recent advancements in self-rewarding language models (SRLMs) (Yuan et al., 2024) have attracted significant attention in the field of LLM alignment due to their ability to efficiently synthesize preference data for iterative preference training. In this approach, a single LLM takes on two roles: the policy model and the reward model. Given unlabeled prompt data, the LLM first acts as the policy model, generating several response candidates. It then functions as the reward model, scoring and ranking these responses. These ranked responses serve as preference pairs to train the LLM using DPO, significantly reducing the reliance on human-annotated data. This process can be iteratively repeated to further enhance performance. However, SRLMs still face challenges in generating reliable and accurate rewards for annotating preference pairs, which is crucial for ensuring the quality of preference data and the alignment performance of LLMs.
To address these challenges, we first formulate a generalized iterative preference fine-tuning framework to analyze the self-rewarding training, where this framework can also be adapted to other iterative preference tuning methods. Through this theoretical framework, we identify that the rewarding bias issue in SRLMs arises from overconfident preference labeling, which forces the model to differentiate between responses of similar quality. For example, both two responses in Figure 1 receive high-quality judgments from human evaluators. However, the SRLM requires the reward model to make a preference judgment, resulting in noisy and unreliable preference labeling. This can negatively impact the model’s preference tuning. Additionally, the iterative training process can further accumulate rewarding bias, diminishing the potential benefits of self-improvement.
@@
Based on insights from our theoretical analysis, we propose Consistency Regularized sElf-rewarding lAnguage Model (CREAM) to mitigate the rewarding bias issue in SRLMs, particularly for widely accessible 7B-size LLMs. The core idea behind CREAM is to avoid forcing the model to be overly confident when distinguishing between responses of similar quality. But how can we determine if the preference labeling is reliable? Outside the self-rewarding scenario, we could employ a pool of external reward models to assist in ranking preferences. When two responses are of similar quality, these external models often yield inconsistent rankings, which signal the confidence level of the preference labeling. In self-rewarding scenarios, however, integrating such external reward models is not feasible. Fortunately, due to the iterative nature of self-rewarding training, we can use the reward model from the previous iteration to rank preferences and then compare these rankings with those produced by the current model. This comparison provides an estimate of such consistency rate. By leveraging this consistency rate, we can regularize the preference training to prevent the model from learning from unreliable preference data, thereby mitigating the rewarding bias issue in SRLMs.
In summary, we first formulate a generalized iterative preference fine-tuning framework to analyze the rewarding bias issue in SRLMs. From the insights of theoretical analysis, we propose CREAM as the primary contribution of this paper. CREAM leverages the rewarding consistency across different iterations for regularized preference training, which can effectively mitigate the rewarding bias issue in SRLMs. Empirical results on a series of natural language benchmarks validate the effectiveness of CREAM in mitigating the rewarding bias issue and enhancing the alignment performance of LLMs.
Notations. Vectors are denoted by lowercase boldface letters, such as $\mathbf{x}$, and matrices by uppercase boldface letters, such as $\mathbf{A}$. For any positive integer $k$, the set ${1,2,\dots,k}$ is denoted by $[k]$. Sets are denoted by calligraphic uppercase letters, such as $\mathcal{D}$, with the cardinality of the set represented as $|\mathcal{D}|$. Without ambiguity, we denote $\pi_{\bm{\theta}}$ as the language model parameterized by $\bm{\theta}$, $\mathbf{x}$ as the input prompt, and $\mathbf{y}$ as the output response from the language model. All other notations are defined prior to their first usage. We denote $\operatorname{\mathds{1}}[\cdot]$ as the indicator function.
@@ 2 Related Works
This paper primarily focuses on addressing the rewarding bias issue in self-rewarding language models (SRLMs) (Yuan et al., 2024), a type of self-improvement method for LLM alignment. In this section, we provide an overview of recent advancements in LLM alignment and delve into a detailed discussion of SRLMs.
@@
LLM Alignment. Alignment is at the core of LLM research and applications, with the goal of ensuring that LLMs adhere to human values and preferences. RLHF established the foundational alignment training paradigm (Leike et al., 2018; Ziegler et al., 2019; Ouyang et al., 2022a), where it leverages human preference feedback to train a reward model, and then use this reward model to guide the LLM via reinforcement learning algorithms (Schulman et al., 2017). Recent efforts have been made to develop direct alignment methods (Rafailov et al., 2023; Dong et al., 2023; Azar et al., 2023; Ethayarajh et al., 2024; Meng et al., 2024; Hong et al., 2024; Zhou et al., 2024a), in order to reduce the costs and complexity of RLHF and make it more efficient and accessible. Representatively, DPO (Rafailov et al., 2023) as a representative direct alignment method, optimizes the LLM with annotated preference pairs, eliminating the need of training an additional reward model. However, most RLHF and direct alignment methods heavily rely on human-annotated preference data, where the data collection commonly involves human to distinguish the “good” responses from the “bad” ones, which is time-consuming and labor-intensive (Ouyang et al., 2022a; Bai et al., 2022). Thus, synthesizing preference data with minimal human efforts has become a valuable research direction.
@@
Self-Rewarding Language Model. SRLM (Yuan et al., 2024) has emerged as a promising approach to address the challenge of preference data synthesis in a self-improvement manner. This method leverages the LLM itself to act as both the policy model and the reward model. The policy model can generate response candidates for unlabeled prompts, while the reward model uses LLM-as-A-Judge (Zheng et al., 2023; Bai et al., 2023; Dubois et al., 2024) prompting to reward and rank these responses based on their quality. The ranked responses are then used as preference pairs to train the LLM via DPO (Rafailov et al., 2023). And this process can be iteratively repeated to improve the alignment performance without human intervention. However, having the same LLM serve as both the policy and reward model, without any regularization, presents challenges in guaranteeing accurate rewards. This can lead to accumulated bias and noisy preference data, which ultimately harms the training. Other similar self-improvement methods (Huang et al., 2022; Zelikman et al., 2022; Chen et al., 2024; Guo et al., 2024b; Zhou et al., 2024b; Wang et al., 2024) often either use the ground truth response to avoid annotation bias, or introduce an additional reward model to reduce the noise in annotations. In contrast, our work neither requires labeled data nor relies on external LLMs. Instead, we propose to use the rewarding consistency to mitigate the rewarding bias in SRLMs.
@@ 3 Methodology
In this section, we first formulate the generalized iterative preference fine-tuning framework for self-rewarding, RL with AI feedback, and other iterative preference fine-tuning methods. Next, we introduce the motivation behind the proposed consistency regularized self-rewarding method. Finally, we present the practical implementation algorithm of CREAM in detail.
@@ 3.1 Generalized Iterative Preference Fine-Tuning Framework
We assume that we can access the dataset with response $\mathcal{D}_{\text{S}}$ and the prompt dataset without response $\mathcal{D}_{\text{U}}$. The objective is to iteratively minimize the following loss with respect to the neural network parameter $\bm{\theta}$ and a label function $z$ as
@@
Different from existing methods, the proposed two-step optimization method directly uses the intrinsic reward model to generate the preference data. This approach is particularly feasible for smaller LLMs, which lack the capacity to effectively use LLM-as-a-Judge prompts (Zheng et al., 2023) for rewarding and ranking. We note that the proposed two-step method is similar to the Expectation-Maximization algorithm and self-training paradigm (Zou et al., 2019). This similarity is supported by the following theorem, which suggests the convergence of the proposed two-step algorithm.
@@ 3.2 Consistency Regularized Self-Rewarding
The generalized framework presented in Eq. (3.1) assumes the human feedback or GPT-4 are all reliable so that the preference labeling function $z$ is trustworthy. However, for SRLMs, the accuracy of preference labeling is not always guaranteed. Therefore, treating all selected preference labels as “ground truth” by encoding them as hard labels can lead to overconfident mistakes, potentially propagating biases and inaccuracies from the LLMs. Taking Figure 1 as an example, both the two responses $\mathbf{y}$ and $\mathbf{y}^{\prime}$ are judged by humans to be of high quality. Forcing the model to be overly confident in distinguishing between these two responses $\{\mathbf{y},\mathbf{y}^{\prime}\}$ with similar quality can negatively impact the performance of SRLMs during training.
This rewarding bias issue motivates us to mitigate such ambiguity by introducing a consistency-regularized self-rewarding language model, CREAM. Specifically, for a pair of responses with very similar quality, their oracle reward scores should ideally be very close to each other. Particularly, when multiple reward models are available, it is likely that some models will rank one response as superior, while others may rank the opposite response as better, resulting in high ranking inconsistency (i.e., low ranking consistency) among these models. Based on this, CREAM aims to prevent the model from learning from preference pairs with low consistency. Instead, it focuses solely on preference pairs with high consistency across different reward models, thereby mitigating the rewarding bias issue and stabilizing the learning process to some extent. From the theoretical perceptive, we can introduce a regularization term to Eq. (3.1) as
@@
where the regularization term $\mathcal{L}_{\text{Reg}}(\bm{\theta};\mathbf{y},\mathbf{y}^{\prime},\mathbf{x})$ prevents the model $\pi_{\bm{\theta}}$ from overconfident in distinguishing the preference of $\{\mathbf{y},\mathbf{y}^{\prime}\}$ with similar quality, which is quantified in the following lemma.
@@
is equivariant with minimizing
@@
where the $1-z$ reverses the preference order of $z(\mathbf{y},\mathbf{y}^{\prime},\mathbf{x})$ and $\mathcal{C}_{\lambda}=(1+\lambda)/(1+2\lambda)$.
Theorem 3.3 suggests that instead of calculating the regularization term $\mathcal{L}_{\text{Reg}}$, we can use the soft-labeled DPO to train Eq. (3.7). In particular, when $\lambda=0$, $\mathcal{C}_{\lambda}=0$ and Eq. (3.7) degenerates to Eq. (3.1). This represents the case where the preference label $z$ is trustworthy from human or some oracle reward models (e.g., GPT-4). In other words, $\lambda$ represents the confidence of the label function $z$. Specially, since in our two-step optimization paradigm, the label function $z$ is directly derived from the previous model $\pi_{\bm{\theta}_{t}}$, we can measure the performance of $\pi_{\bm{\theta}_{t}}$ using the consistency between model $\bm{\theta}_{t}$ and the baseline reward model (e.g., external reward model) $\bm{\theta}_{t}^{\prime}$, defined by

### v3

@@ Abstract
Recent self-rewarding large language models (LLM) have successfully applied LLM-as-a-Judge to iteratively improve the alignment performance without the need of human annotations for preference data. These methods commonly utilize the same LLM to act as both the policy model (which generates responses) and the reward model (which scores and ranks those responses). The ranked responses are then used as preference pairs to train the LLM via direct alignment technologies (e.g. DPO). However, it is noteworthy that throughout this process, there is no guarantee of accuracy in the rewarding and ranking, which is critical for ensuring accurate rewards and high-quality preference data. Empirical results from relatively small LLMs (e.g., 7B parameters) also indicate that improvements from self-rewarding may diminish after several iterations in certain situations, which we hypothesize is due to accumulated bias in the reward system. This bias can lead to unreliable preference data for training the LLM. To address this issue, we first formulate and analyze the generalized iterative preference fine-tuning framework for self-rewarding language model. We then introduce the regularization to this generalized framework to mitigate the overconfident preference labeling in the self-rewarding process. Based on this theoretical insight, we propose a Consistency Regularized sElf-rewarding lAnguage Model (CREAM) that leverages the consistency of rewards across different iterations to regularize the self-rewarding training, helping the model to learn from more reliable preference data. With this explicit regularization, our empirical results demonstrate the superiority of CREAM in improving both reward consistency and alignment performance. The code is publicly available at https://github.com/Raibows/CREAM.
@@ 1 Introduction
Large language models (LLMs) have shown impressive capabilities across various tasks, including natural language understanding and generation (Radford et al., 2019). At the same time, LLMs also face alignment challenges such as generating hallucinations and harmful outputs (Ji et al., 2023). To address these issues, a series of research works has explored preference learning methods such as Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017) and direct alignment techniques such as Direct Preference Optimization (DPO) (Rafailov et al., 2023) to align the LLMs with human values and preferences. These alignment methods often require a large number of preference pairs which are indispensable in both RLHF and direct alignment training. However, collecting human-annotated preference pairs is time-consuming and labor-intensive, which seriously limits the scalability and efficiency of these alignment methods.
Recent advancements in self-rewarding language models (SRLMs) (Yuan et al., 2024) have attracted increasing attention in the field of LLM alignment, which can efficiently synthesize preference data for iterative preference training. In this method, a single LLM is required to act as two roles, the policy model and the reward model. Given unlabeled prompt data, the LLM first acts as the policy model generating several response candidates. Then, the same LLM acts as the reward model, scoring and ranking these responses. These ranked responses are used as preference pairs to train the LLM with DPO, significantly reducing the reliance on human-annotated data. The above steps can be iteratively repeated to further enhance the performance. However, SRLMs still face challenges in generating reliable and accurate rewards for annotating the preference pairs, which is critical for ensuring both the quality of preference data and the alignment performance of LLMs.
To address these challenges, we first formulate a generalized iterative preference fine-tuning framework to analyze the self-rewarding training, where this framework can also be adapted to other iterative preference tuning methods. Through this theoretical framework, we find that the rewarding bias issue in SRLMs arises from the overconfident preference labeling, which enforces the model to distinguish between responses of similar quality. For example, both responses in Figure 1 have high quality judgments from the human. The SRLM enforces the reward model to make a preference judgment, resulting in noisy and unreliable preference labeling. This can lead to negative impacts on preference tuning the model. Additionally, the iterative training manner can also accumulate the rewarding bias, further diminishing the benefits of self-improvement.
@@
From the insights of theoretical analysis, we propose Consistency Regularized sElf-rewarding lAnguage Model (CREAM) to mitigate the rewarding bias issue in SRLMs, particularly for broadly accessible 7B-size LLMs. The core idea behind CREAM is that we should not force the model to be overly confident in distinguishing between responses of similar quality. But how to tell the preference labeling is reliable or not? Out of the self-rewarding scenario, we may employ a pool of external reward models to assist in ranking preferences. When two responses are of similar quality, these external models often produce inconsistent rankings. This inconsistency serves as a signal to indicate the level of confidence in the preference labeling. In self-rewarding scenarios, however, integrating such external reward models is not feasible. Fortunately, due to the iterative nature of self-rewarding training, we can use the reward model from the previous iteration to rank preferences and then compare these rankings with those produced by the current model. This comparison provides an estimate of such consistency rate. With this consistency rate, we can regularize the preference training to prevent the model from learning unreliable preference data, thereby mitigating the rewarding bias issue in SRLMs.
In summary, we first formulate a generalized iterative preference fine-tuning framework to analyze the rewarding bias issue in SRLMs. From the insights of theoretical analysis, we propose CREAM as the primary contribution of this paper. CREAM leverages the consistency of rewarding across different iterations for regularized preference training, which can effectively mitigate the rewarding bias issue in SRLMs. Empirical results on a series of natural language benchmarks validate the effectiveness of CREAM in mitigating the rewarding bias issue and enhancing the alignment performance of LLMs.
Notations. Vectors are denoted by lowercase boldface letters, such as $\mathbf{x}$, and matrices by uppercase boldface letters, such as $\mathbf{A}$. For any positive integer $k$, the set ${1,2,\dots,k}$ is denoted by $[k]$. Other general sets are denoted by calligraphic uppercase letters, such as $\mathcal{D}$, with the cardinality of the set represented as $|\mathcal{D}|$. Without ambiguity, we denote $\pi_{\bm{\theta}}$ as the language model parameterized by $\bm{\theta}$, $\mathbf{x}$ as the input prompt, and $\mathbf{y}$ as the output response from the language model. All other notations are defined prior to their first usage. We denote $\operatorname{\mathds{1}}[\cdot]$ as the indicator function.
@@ 2 Related Works
LLM Alignment. Alignment lies at the core of LLM research and applications, aiming to ensure that LLMs adhere to human values and preferences. RLHF establishes the foundational alignment training paradigm (Leike et al., 2018; Ziegler et al., 2019; Ouyang et al., 2022a), where it leverages human preference feedback to train a reward model, and then uses this reward model to guide the LLM via reinforcement learning algorithms (Schulman et al., 2017). Recent efforts have focuses on developing direct alignment methods (Rafailov et al., 2023; Dong et al., 2023; Azar et al., 2023; Ethayarajh et al., 2024; Meng et al., 2024; Hong et al., 2024), in order to reduce the costs and complexity of RLHF and make it more efficient and accessible. DPO (Rafailov et al., 2023), as a representative direct alignment method, optimizes the LLM with annotated preference pairs, eliminating the need for training an additional reward model. However, most RLHF and direct alignment methods heavily rely on human-annotated preference data, where the data collection commonly involves human distinguishing the “good” responses from the “bad” ones, which is time-consuming and labor-intensive (Ouyang et al., 2022a; Bai et al., 2022). Thus, synthesizing preference data with minimal human effort has become a valuable research direction.
@@
Self-Rewarding Language Model. SRLM (Yuan et al., 2024) has emerged as a promising approach to address the challenge of preference data synthesis in a self-improvement manner. This method leverages the LLM itself to act as both the policy model and the reward model. The policy model can generate response candidates for unlabeled prompts, while the reward model uses LLM-as-A-Judge (Zheng et al., 2023; Bai et al., 2023; Dubois et al., 2024) prompting to reward and rank these responses based on their quality. The ranked responses are then used as preference pairs to train the LLM via DPO (Rafailov et al., 2023). And this process can be iteratively repeated to improve the alignment performance without human intervention. However, having the same LLM serve as both the policy and the reward model, without any regularization, presents challenges in guaranteeing accurate rewards. This can lead to bias accumulation and noisy preference data, which ultimately harms the training. Other similar self-improvement methods (Huang et al., 2022; Zelikman et al., 2022; Chen et al., 2024; Guo et al., 2024b; Zhou et al., 2024) often either use the ground truth response to avoid annotation bias, or introduce an additional reward model to reduce the noise in annotations. In contrast, our work neither requires labeled data nor relies on external LLMs. Instead, we propose to use the consistency of rewarding to mitigate the rewarding bias in SRLMs.
Reward Hacking. In both RLHF and SRLM scenarios, the reward model plays a crucial role in training LLMs (Ouyang et al., 2022b; Anwar et al., 2024; Yuan et al., 2024; Fisch et al., 2024). For RLHF, reward hacking is a phenomenon where models exploit flaws or biases in reward models to maximize scores without aligning with the intended goals (Anwar et al., 2024). To mitigate this issue, various ensemble rewarding methods (Coste et al., 2023; Eisenstein et al., 2023; Ramé et al., 2024; Zhang et al., 2024) such as ensemble-based conservative optimization (Coste et al., 2023) and averaging rewards in the weight space (Ramé et al., 2024) have been proposed to improve reliability and robustness. However, these works mainly focused on estimating the rewards, while CREAM in the self-rewarding scenario only uses rewards for comparing the responses to annotate preference pairs instead of maximizing the rewards. Besides, CREAM uses regularization instead of conservative value estimation to mitigate the rewarding bias issue. Though applying ensemble rewarding methods such as robust preference optimization (Fisch et al., 2024) which proposes reward distillation and pessimistic ensemble is feasible, CREAM does not rely on ensemble rewarding which reduces the computational cost in SRLMs. Additionally, CREAM leverages the iterative nature of SRLM to estimate the consistency between iterations to regularize the preference pair, which is more suitable for iterative preference fine-tuning. Further discussion is in Appendix C.4.
@@ 3 Methodology
In this section, we first formulate the generalized iterative preference fine-tuning framework for self-rewarding, RL with AI feedback, and other iterative preference tuning methods. Next, we introduce the motivation behind the proposed consistency regularized self-rewarding method. Finally, we present the practical implementation algorithm of CREAM in detail.
@@ 3.1 Generalized Iterative Preference Fine-Tuning Framework
We assume that we can access to the dataset with response $\mathcal{D}_{\text{S}}$ and the prompt dataset without response $\mathcal{D}_{\text{U}}$. The objective is to iteratively minimize the following loss with respect to the neural network parameter $\bm{\theta}$ and a label function $z$ as
@@
Different from existing methods, the proposed two-step optimization method directly uses the intrinsic reward model to generate the preference data. This approach is particularly feasible for small LLMs, which lack the capacity to effectively use LLM-as-a-Judge prompts (Zheng et al., 2023) for rewarding and ranking. We note that the proposed two-step method is similar to the Expectation-Maximization algorithm and self-training paradigm (Zou et al., 2019). This similarity is supported by the following theorem, which suggests the convergence of the proposed two-step algorithm.
@@ 3.2 Consistency Regularized Self-Rewarding
The generalized framework presented in Eq. (3.1) assumes the human feedback or GPT-4 are all reliable so that the preference labeling function $z$ is trustworthy. However, for SRLMs, the accuracy of preference labeling is not always guaranteed. Therefore, treating all selected preference labels as “ground truth” by encoding them as hard labels can lead to overconfident mistakes, potentially propagating biases and inaccuracies from the LLMs. Taking Figure 1 as an example, both the two responses $\mathbf{y}$ and $\mathbf{y}^{\prime}$ are judged by humans to be of high quality. Forcing the model to be overly confident in distinguishing between these two responses $\{\mathbf{y},\mathbf{y}^{\prime}\}$ of similar quality can negatively impact the performance of SRLMs during training.
This rewarding bias issue motivates us to mitigate such ambiguity by introducing a consistency-regularized self-rewarding language model, CREAM. Specifically, for a pair of responses with very similar quality, their oracle reward scores should ideally be very close to each other. Particularly, when multiple reward models are available, it is likely that some models will rank one response as superior, while others may rank the opposite response as better, resulting in high ranking inconsistency (i.e., low ranking consistency) among these models. Based on this, CREAM aims to prevent the model from learning from preference pairs with low consistency. Instead, it focuses solely on preference pairs with high consistency across different reward models, thereby mitigating the rewarding bias issue and stabilize the learning process to some extent. From the theoretical perceptive, we can introduce a regularization term to Eq. (3.1) as
@@
where the regularization term $\mathcal{L}_{\text{Reg}}(\bm{\theta};\mathbf{y},\mathbf{y}^{\prime},\mathbf{x})$ prevents the model $\pi_{\bm{\theta}}$ from overconfidence in distinguishing the preference of $\{\mathbf{y},\mathbf{y}^{\prime}\}$ of similar quality, which is quantified in the following lemma.
@@
is equivalent with minimizing
@@
where the $1-z$ reverses the preference order of $z(\mathbf{y},\mathbf{y}^{\prime},\mathbf{x})$ and $\mathcal{C}_{\lambda}=(1+\lambda)/(1+2\lambda)$. We choose to reverse the preference order if there is evidence that the annotated preference data is reversed. And the final form can also be viewed as the label smoothing. Details are in Appendix C.2.
Theorem 3.3 suggests that instead of calculating the regularization term $\mathcal{L}_{\text{Reg}}$, we can use the soft-labeled DPO to train Eq. (3.7). In particular, when $\lambda=0$, $\mathcal{C}_{\lambda}=0$ and Eq. (3.7) degenerates to Eq. (3.1). This represents the case where the preference label $z$ is trustworthy from human or some oracle reward models (e.g., GPT-4). In other words, $\lambda$ represents the confidence of the label function $z$. Specially, since in our two-step optimization paradigm, the label function $z$ is directly derived from the previous model $\pi_{\bm{\theta}_{t}}$, we can measure the performance of $\pi_{\bm{\theta}_{t}}$ using the consistency between model $\bm{\theta}_{t}$ and the baseline model (e.g., external reward model) $\bm{\theta}_{t}^{\prime}$, defined by